Timed Teit Wamings appear when half the time Multiple Alles Not This test does not aon Remaining Time: 15 minubes 15secends Quiestion 3 An aqueous soluton was found to contar, a mn s or su d a nr nitrate anNormu tras1894 ghrein, ww owe a the molarity concentradion of the soluse in the soluzion mL Cal ?ge 3730 M 3.73 M 0
Answer #1

molar mass of Zn(NO3)2=189.4 g/mole

moles of Zn (NO3)2= mass of Zn(NO3)2/ molar mass = 25.38/189.4 =0.134

volume of solution = 500ml, 1L= 1000ml

concentration of solution in liters= 500/1000=0.5L

concentration of Zn(NO3)2= moles of Zn (NO3)2/volume in liters =0.134/0.5= 0.268M ( D is corect answder)